OpenStudy (anonymous):

What is the perimeter, in square centimeters, of a rectangle that has a length of 4 centimeters and a width of 15 millimeters? 38 cm 19 cm 5.5 cm 11 cm

OpenStudy (kash_thesmartguy):

2(l+w) is the formula for a rectangle perimeter. So 15 mm=1.5 cm. Then, 2(4+1.5)= Your answer.

OpenStudy (kash_thesmartguy):

\[2(4+1.5)=11ft ^{2}\] 11 ft2 is you answer

OpenStudy (anonymous):

To get from one town to another, Pierre must travel 15 kilometers west and 8 kilometers north. What is the direct distance between the two towns? 23 km 13 km 17 km 7 km

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

give metal plz anyone

OpenStudy (kash_thesmartguy):

15+8=? (? is your answer) It's 23, so A.

OpenStudy (anonymous):

Which of the following sets of numbers is a Pythagorean triple? 42, 40, 82 42, 40, 58 41, 41, 58 40, 41, 57